Water of density 4kg/ m^3 and ice of density 2kg/m^3 are mixed together. If their masses are equal then the density of mixtures is

Grade:8

3 Answers

Arun
4 years ago
 
If two loquid lf equal mass and different densities are mixed then
Density = 2 *d1 *d2/(d1 +d2)
 = 8/3 kg/m³
Vikas TU
4 years ago
Dear student 
See the explanation here. 
The masses of two liquids are equal, let it be m.
Let the relative densities of water and liquid be ρ1  and ρ2 respectively.
The volume of the two be V1 and V2, of water and liquid respectively.
The volume of the mixture would be, V = V1 + V2  (1)
also, volume = mass/density 
thus, 
2m/ρ (V) = m/ρ1 (V 1 ) + m/ρ2 (V 2 )
here ρ1 = 1, ρ2 = 2 and ρ is the relative density of the mixture.
now,
2/ρ = 1/ρ1 + 1/ρ2
by substituting the values, we 
ρ/2 = 2/3
or, the relative density of the combined liquid will be, ρ= 4/3
Vikas TU
4 years ago
Dear student 
The ans of the above question will be 8/3 , it is by mistake written as 4/3. 
Hope you will understand the explanation.
Good Luck